# Local algebraicity and localization of the Bergman kernel on Stein spaces with finite type boundaries

### v1

@@ 1. Introduction
The purpose of this paper is two-fold. The first objective, and the original motivation for the work, is to generalize results from [EXX23] regarding algebraic Bergman kernels on domains in $\mathbb{C}^{2}$ to domains in 2-dimensional Stein spaces. The second, needed to resolve the first, is to establish a result on localization of the Bergman kernel on domains with smooth finite type boundaries in complex manifolds. The second objective is of independent interest and generalizes the classical work of Fefferman [Fef74] for strictly pseudoconvex domains in $\mathbb{C}^{n}$, as well as other work in more general situations (see below). The proof of the localization result follows the general approach in [Fef74, HL20] but requires some new technology, including the introduction of a new negative Sobolev norm.
In complex analysis and geometry, the Bergman kernel and metric have been fundamental objects for our understanding of the geometry of complex manifolds since the classical work of Bergman and [Ber33], [Ber35] and Kobayashi [Kob59]. For example, the resolution of Cheng’s Conjecture [FW97], [NS06], [HX21] characterizes the unit ball $\mathbb{B}^{n}\subset\mathbb{C}^{n}$, up to biholomorphisms, as the only smoothly bounded strictly pseducoconvex domain in $\mathbb{C}^{n}$ with a Bergman metric that is also Kähler-Einstein. Further developments on generalizations of Cheng’s Conjecture can be found in, e.g., [HL20], [EXX22], [GS22] and [SX23]. In a different direction, together with H. Xu the first and third authors [EXX24, EXX23] studied bounded pseudoconvex domains $G$ in $\mathbb{C}^{2}$ whose Bergman kernels are algebraic. In [EXX24], it is proved that if the boundary of $G$ is strongly pseudoconvex and its Bergman kernel is algebraic, then there is an algebraic biholomorphism from $G$ to $\mathbb{B}^{2}$. This conclusion fails if the strong pseudoconvexity of $\partial G$ is dropped in view of calculations due to D’Angelo [D’A78] showing that complex ellipsoids of the form
@@
We make several remarks. First, we shall refer to a Bergman kernel form $K=K_{\Omega}(z,\bar{z})$ (on the diagonal) such that there is a local coordinate chart at $\xi\in\partial\Omega$ in which $K$ is algebraic with algebraic degree $\leq d$ as locally algebraic at $\xi$ of degree $\leq d$. It is (essentially) proved in [EXX24] that if the boundary $\partial\Omega$ is strongly pseudoconvex and $K$ is locally algebraic, then the log-term in Fefferman’s asymptotic expansion of $K$ vanishes to infinite order at $\partial\Omega$; thus, having a locally algebraic Bergman kernel form is a strong (non-generic) condition. Second, we mention that the degree estimate in Theorem 1 (i) is sharp in view of the result by D’Angelo (loc. cit.) on the Bergman kernels of the complex ellipsoids mentioned above. Third, the converse in Theorem 1 (ii), namely that ball quotients have locally rational Bergman kernels was shown in [EXX22] (see also [HL20]). Fourth, by an example in [EXX24, Section 6], the conclusion in Theorem 1 (ii) fails in higher dimensions; The example in [EXX24, Section 6] is a domain in a 3-dimensional Stein space (with a single isolated singularity) that is not biholomorphic to a ball quotient, but whose Bergman kernel can be shown to be locally rational at every boundary point (a minor detail is left to the reader). We finally note that, as pointed out in [EXX23], one cannot expect to have an upper bound on the algebraic degree of Bergman kernel in terms of finite type of the boundary. This is because the latter is a biholomorphic invariant while the algebraic degree is not, and can be made arbitrarily large by biholomorphic transformations.
@@
Theorem 2 follows immediately from Lemma 2.1 and Theorem 3 below. For the proof, we first follow the framework of [Fef74, HL20]. We need to introduce some new ideas for this approach to work in the finite type case. In particular, we introduce a new Sobolev-type norm for negative indices, which is used in our proof.
@@ Acknowledgement
The authors are grateful to Ioan Bejenaru, Harold Boas, Xiaoshan Li, Mei-Chi Shaw, Ziming Shi, and Emil Straube for important insights regarding $\bar{\partial}$-Neumann problem, techniques in PDEs, and harmonic analysis.
@@ 2. Localization of the Bergman kernel
In this section we prove a localization result for the Bergman kernel of a complex manifold with pseudoconvex and smooth, finite type boundary. There is such a localization result in [Eng01] for finite type, pseudoconvex domains in $\mathbb{C}^{n}$. However the approach in that work cannot be directly carried over to domains in complex manifolds and for our application purpose, we need to extend the result to the manifold case. Huang-Li already established a localization result for the Bergman kernel of complex manifolds with strongly pseudoconvex boundary in [HL20]. For our proof, we shall follow the basic approach in [HL20], but make necessary changes and adjustments for the context of finite type. We recall that there are two standard notions of finite type, one due to Hörmander, Kohn, Bloom–Graham and the other due to D’Angelo. These notions coincide for a real hypersurface in $\mathbb{C}^{2}$ or, more generally, in a complex manifold of dimension two. Our localization result here is for $n\geq 2$ dimensions and we will consider finite D’Angelo 1-type, as it gives us the necessary and sufficient condition for existence of subelliptic estimates of the $\bar{\partial}$-Neumann problem [Cat83, Cat87a] (see also [CD10]). The precise definition of D’Angelo 1-type (henceforth, simply referred to as ”finite type”) is not needed for our purposes, only the subelliptic estimates, and the interested reader is referred to [D’A93] (see also [BER99]) for the definition.
@@
If we compare these two notions of negative Sobolev norms, we easily see that $|\!|{}g|\!|{}_{-s}\leq|\!|{}g|\!|{}^{\dagger}_{-s}$. We define $H^{\dagger}_{-s}(M)$ to be the space of distributions acting on the space $\Omega^{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*})$ with finite ||.$|\!|{}.|\!|{}^{\dagger}_{-s}$ norm. In particular, $H^{\dagger}_{-s}(M)$ includes every $f\in L^{2}_{p,q}(M)$ that acts naturally on $\Omega^{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*})$ by the $L^{2}$ inner product defined above.
@@
for some sufficiently small $\delta>0$, where $B(p,\epsilon)=\left\{q\in U:|z(q)|=\sqrt{\sum_{j=1}^{n}|z_{j}(q)|^{2}}<\epsilon\right\}$.
@@
The proof uses the general theory and framework of the $\bar{\partial}$-Neumann problem and follows the approach in the proof of Proposition 3.1 in [HL20]. A reader not familiar with the $\bar{\partial}$-Neumann problem is referred to, e.g., [FK72] for an introduction. We split the proof into two steps.
Step 1: Fix a holomorphic chart $(U,w)$ at $p$, and as before, we write $d\bar{w}|_{w}=d\bar{w}_{1}\wedge\cdots\wedge d\bar{w}_{n}|_{w}$ for $w\in U$. Fix some $w\in B(p,\delta)\cap M$ and set
@@
where $\chi_{D}$ is the characteristics function of $D$. We write $f_{w}(z)\coloneqq\tilde{f}_{w}(z)\wedge d\bar{w}|_{w}$ and $\tilde{g}_{w}(z)\coloneqq\bar{\partial}\tilde{f}_{w}(z)$. Thus, $\tilde{f}_{w}(z)$ is an $L^{2}$ integrable (with respect to $z$ coordinate) $(n,0)$ form on $M$. By the reproducing property of Bergman kernel, $\tilde{f}_{w}(\cdot)\perp A^{2}(M)$ and $\tilde{g}_{w}$ is a $(n,1)$ form in the sense of distributions. We note that $\tilde{g}_{w}\in H^{\dagger}_{-1}(M)$, $supp(\tilde{g}_{w}(z))\subset\partial D\setminus\partial M$, and $(\tilde{g}_{w},h)=(\bar{\partial}\tilde{f}_{w},h)\coloneqq(\tilde{f}_{w}(z),\bar{\partial}^{*}h)$ for $h\in\Omega^{n,0}(\overline{M})\cap Dom(\bar{\partial}^{*})$. Pick a sequence of $(n,0)$ forms $\{\tilde{f}_{w,j}\}_{j\in\mathbb{N}}$ on $M$ which are smooth in $\overline{M}$ (we in fact can take them to be compactly supported in $M$) such that $\tilde{f}_{w,j}\to\tilde{f}_{w}$ in $L^{2}_{n,0}(M)$. Set $\tilde{g}_{w,j}(z)\coloneqq\bar{\partial}\tilde{f}_{w,j}.$ Since $supp(\tilde{g}_{w}(z))\subset\partial D\setminus\partial M$, we can assume that $supp(\tilde{g}_{w,j}(z))$ is contained in a $2^{-j}$-neighborhood (with respect to some complete metric on $M$ that may be different from $g^{\prime}|_{M}$) of $supp(\tilde{g}_{w}(z))$ in $M$ for $j$ large enough. Moreover, we have
@@
Let $N:L^{2}_{n,1}(M)\to L^{2}_{n,1}(M)$ be the $\bar{\partial}$-Neumann operator. Let $p\in\partial M$ have finite type $m$. Then the following pseudolocal estimate holds for $g=\bar{\partial}f$ such that $f$ a smooth $(n,0)$ form $f$ and $g$ has compact support in $U\cap\overline{M}$:
@@
for any $\xi,\xi_{1}\in C^{\infty}_{c}(B(p,\frac{3}{2}\delta))$ that satisfy $\xi_{1}|_{W}\equiv 1$ and $\xi|_{B(p,\delta)}\equiv 1$, where $W\subset B(p,\frac{3}{2}\delta)$ is a small neighborhood of $supp(\xi)$. The constants $\{C_{s}\}$ depend on $s$ and $\xi$, but not on $W$ and $\xi_{1}$. In the two dimensional case, $\lambda$ can be taken as $2^{-m}$ (by [Koh72]) and $\leq 1/m$ in higher dimensions (by [Cat87b, Theorem 3.3]).
@@
These pseudolocal estimates essentially follow from the subelliptic estimates of the $\bar{\partial}$-Neumann operator, as mentioned in [FK72]. In two dimensions, subelliptic estimates are proven in [Koh72]. In higher dimensions, subelliptic estimates for D’Angelo finite type domains in $\mathbb{C}^{n}$ were proved by Catlin (see [CD10], [Cat87b]); these are also valid in our manifold setting because of their local nature. In the following we give a proof of how these pseudolocal estimates follow from the subelliptic ones for manifolds. We also note that the proof of existence of the operator $N$ for $(n,1)$ forms is similar to that for $(0,1)$ forms.
@@
To apply Lemma B.1, we let $\xi,\xi_{1}$ be as in Claim 2.1. Let $d:=dist(supp(\xi),supp(1-\xi_{1}))$. Note that by a suitable scaling of coordinates, we may assume $d\leq 1$. Furthermore, we take $\epsilon=\lambda$, $r=s+\lambda$, $m=1$, $m^{\prime}=1+2\lambda$, $L=N$, $f=\tilde{g}_{w,j}$, $R(\cdot)=|\!|{}\cdot|\!|{}_{s}$, $T=id$ and $G(\cdot)=|\!|{}\cdot|\!|{}^{\dagger}_{-1}$ in Lemma B.1. By (2.7), we see that the inequality (B.2) holds. Note that to apply Lemma B.1, we also need $N$ to be a linear operator from $L^{2}(M)$ to itself, and that it satisfies (B.1) with $\Omega$ replaced by $M$. This is verified in Proposition C.1 in Appendix C. Applying Lemma B.1 then gives us (2.6). ∎
@@
Let $\sigma>0$ be small and $\rho\in C^{\infty}(M^{\prime})$ be a real-valued function so that $\rho$ is compactly supported in a $2\sigma$ neighborhood of $\overline{\partial D\setminus\partial M}$ in $M$, with respect to the Euclidean metric on $U$, which we denote by $V_{2\sigma}$ and is identically one in $V_{\sigma}$. Write $K_{D}(z,w)=\tilde{K}_{D}(z,\bar{w})\wedge d\bar{w}|_{w}$ for all $w\in B(p,\delta)\cap M$. Fix any $\phi=\sum_{j=1}^{n}\phi_{j}dz_{1}\wedge\cdots\wedge dz_{n}\wedge d\bar{z}_{j}\in\Omega^{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*})$. Since $supp(\tilde{g}_{w})\subset\partial D\setminus\partial M$, by our definition of $\tilde{g}_{w}$ acting on $\Omega^{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*})$, we have $(\tilde{g}_{w},\phi)=(\tilde{g}_{w},\rho\phi)$ and
@@
where $\tilde{K}_{D}(z,\bar{w})=k_{D}(z,\bar{w})dz_{1}\wedge\cdots\wedge dz_{n}$ in local coordinates. Assuming $\sigma,\delta$ are sufficiently small, we observe that $dist(V_{2\sigma},B(p,\delta))>0$. By Theorem 1 in [Bel86], it follows that

### v2

@@ 1. Introduction
The purpose of this paper is two-fold. The first objective, and the original motivation for the work, is to generalize results from [EXX23] regarding algebraic Bergman kernels on domains in $\mathbb{C}^{2}$ to domains in 2-dimensional Stein spaces. The second, needed to resolve the first, is to establish a result on localization of the Bergman kernel on domains with smooth finite type boundaries in complex manifolds. The second objective is of independent interest and generalizes the classical work of Fefferman [Fef74] for strictly pseudoconvex domains in $\mathbb{C}^{n}$, as well as other work in more general situations (see below). The proof of the localization result follows the general approach in [Fef74, HL23] but requires some new technology, including the introduction of a new negative Sobolev norm.
In complex analysis and geometry, the Bergman kernel and metric have been fundamental objects for our understanding of the geometry of complex manifolds since the classical work of Bergman and [Ber33], [Ber35] and Kobayashi [Kob59]. For example, the resolution of Cheng’s Conjecture [FW97], [NS06], [HX21] characterizes the unit ball $\mathbb{B}^{n}\subset\mathbb{C}^{n}$, up to biholomorphisms, as the only smoothly bounded strictly pseducoconvex domain in $\mathbb{C}^{n}$ with a Bergman metric that is also Kähler-Einstein. Further developments on generalizations of Cheng’s Conjecture can be found in, e.g., [HL23], [EXX22], [GS22] and [SX23]. In a different direction, together with H. Xu the first and third authors [EXX24, EXX23] studied bounded pseudoconvex domains $G$ in $\mathbb{C}^{2}$ whose Bergman kernels are algebraic. In [EXX24], it is proved that if the boundary of $G$ is strongly pseudoconvex and its Bergman kernel is algebraic, then there is an algebraic biholomorphism from $G$ to $\mathbb{B}^{2}$. This conclusion fails if the strong pseudoconvexity of $\partial G$ is dropped in view of calculations due to D’Angelo [D’A78] showing that complex ellipsoids of the form
@@
We make several remarks. First, we shall refer to a Bergman kernel form $K=K_{\Omega}(z,\bar{z})$ (on the diagonal) such that there is a local coordinate chart at $\xi\in\partial\Omega$ in which $K$ is algebraic with algebraic degree $\leq d$ as locally algebraic at $\xi$ of degree $\leq d$. It is (essentially) proved in [EXX24] that if the boundary $\partial\Omega$ is strongly pseudoconvex and $K$ is locally algebraic, then the log-term in Fefferman’s asymptotic expansion of $K$ vanishes to infinite order at $\partial\Omega$; thus, having a locally algebraic Bergman kernel form is a strong (non-generic) condition. Second, we mention that the degree estimate in Theorem 1 (i) is sharp in view of the result by D’Angelo (loc. cit.) on the Bergman kernels of the complex ellipsoids mentioned above. Third, the converse in Theorem 1 (ii), namely that ball quotients have locally rational Bergman kernels was shown in [EXX22] (see also [HL23]). Fourth, by an example in [EXX24, Section 6], the conclusion in Theorem 1 (ii) fails in higher dimensions; The example in [EXX24, Section 6] is a domain in a 3-dimensional Stein space (with a single isolated singularity) that is not biholomorphic to a ball quotient, but whose Bergman kernel can be shown to be locally rational at every boundary point (a minor detail is left to the reader). We finally note that, as pointed out in [EXX23], one cannot expect to have an upper bound on the algebraic degree of Bergman kernel in terms of finite type of the boundary. This is because the latter is a biholomorphic invariant while the algebraic degree is not, and can be made arbitrarily large by biholomorphic transformations.
@@
Theorem 2 follows immediately from Lemma 2.1 and Theorem 3 below. For the proof, we first follow the framework of [Fef74, HL23]. We need to introduce some new ideas for this approach to work in the finite type case. In particular, we introduce a new Sobolev-type norm for negative indices, which is used in our proof.
@@ Acknowledgement
The authors are grateful to Ioan Bejenaru, Harold Boas, Xiaoshan Li, Mei-Chi Shaw, Ziming Shi, Liding Yao, and Emil Straube for important insights regarding $\bar{\partial}$-Neumann problem, techniques in PDEs, and harmonic analysis.
@@ 2. Localization of the Bergman kernel
In this section we prove a localization result for the Bergman kernel of a complex manifold with pseudoconvex and smooth, finite type boundary. There is such a localization result in [Eng01] for finite type, pseudoconvex domains in $\mathbb{C}^{n}$. However the approach in that work cannot be directly carried over to domains in complex manifolds and for our application purpose, we need to extend the result to the manifold case. Huang-Li already established a localization result for the Bergman kernel of complex manifolds with strongly pseudoconvex boundary in [HL23]. For our proof, we shall follow the basic approach in [HL23], but make necessary changes and adjustments for the context of finite type. We recall that there are two standard notions of finite type, one due to Hörmander, Kohn, Bloom–Graham and the other due to D’Angelo. These notions coincide for a real hypersurface in $\mathbb{C}^{2}$ or, more generally, in a complex manifold of dimension two. Our localization result here is for $n\geq 2$ dimensions and we will consider finite D’Angelo 1-type, as it gives us the necessary and sufficient condition for existence of subelliptic estimates of the $\bar{\partial}$-Neumann problem [Cat83, Cat87] (see also [CD10]). The precise definition of D’Angelo 1-type (henceforth, simply referred to as “finite type”) is not needed for our purposes, only the subelliptic estimates, and the interested reader is referred to [D’A93] (see also [BER99]) for the definition.
@@
If we compare these two notions of negative Sobolev norms, we easily see that $|\!|{}g|\!|{}_{-s}\leq|\!|{}g|\!|{}^{\dagger}_{-s}$. We define $H^{\dagger}_{-s}(M)$ to be the completion of $\Omega^{p,q}(\overline{M})$ under ||.$|\!|{}.|\!|{}^{\dagger}_{-s}$ norm. In particular, $H^{\dagger}_{-s}(M)$ includes every $f\in L^{2}_{p,q}(M)$.
@@
for some sufficiently small $10<\delta<$ 1, where $B(p,\epsilon)=\left\{q\in U:|z(q)|=\sqrt{\sum_{j=1}^{n}|z_{j}(q)|^{2}}<\epsilon\right\}$.
@@
The proof uses the general theory and framework of the $\bar{\partial}$-Neumann problem and follows the approach in the proof of Proposition 3.1 in [HL23]. A reader not familiar with the $\bar{\partial}$-Neumann problem is referred to, e.g., [FK72] for an introduction. We split the proof into two steps.
Step 1: Fix the holomorphic chart $(U,w)$ at $p$ as specified in Lemma 2.1, and as before, we write $d\bar{w}|_{w}=d\bar{w}_{1}\wedge\cdots\wedge d\bar{w}_{n}|_{w}$ for $w\in U$. Fix some $w\in B(p,\delta)\cap M$ and set
@@
where $\chi_{D}$ is the characteristics function of $D$. We write $f_{w}(z)\coloneqq\tilde{f}_{w}(z)\wedge d\bar{w}|_{w}$ and $\tilde{g}_{w}(z)\coloneqq\bar{\partial}\tilde{f}_{w}(z)$. Thus, $\tilde{f}_{w}(z)$ is an $L^{2}$ integrable (with respect to $z$ coordinate) $(n,0)$ form on $M$. By the reproducing property of Bergman kernel, $\tilde{f}_{w}(\cdot)\perp A^{2}(M)$ and $\tilde{g}_{w}$ is a $(n,1)$ form in the sense of distributions. We note that $\tilde{g}_{w}\in H^{\dagger}_{-1}(M)$, $supp(\tilde{g}_{w}(z))\subset\partial D\setminus\partial M$, and $(\tilde{g}_{w},h)=(\bar{\partial}\tilde{f}_{w},h)\coloneqq(\tilde{f}_{w}(z),\bar{\partial}^{*}h)$ for $h\in\Omega^{n,0}(\overline{M})\cap Dom(\bar{\partial}^{*})$.
Let us write $\tilde{f}_{w}(z)\coloneqq\breve{f}_{w}(z)dz|_{z}$. If $M$ is defined by a smooth defining function $\{\rho<0\}$, let $\hat{N}$ be the unit outward normal at $p$, i.e., $\vec{N}/||\vec{N}||$ where $\vec{N}=(\frac{\partial\rho}{\partial\bar{z}_{1}}(p),\ldots,\frac{\partial\rho}{\partial\bar{z}_{n}}(p))$. Let us choose $\psi(z)$ as the signed-Euclidean distance function to $\Gamma=\overline{\partial D\setminus\partial M}$ in a small neighborhood $V$ of $\Gamma$ and we extend this smoothly to $M^{\prime}$ such that $\psi<0$ in $D$, $\psi>0$ in $M\setminus\bar{D}$, and $|\psi|\geq\epsilon_{0}$ outside this neighborhood $V$, for some $\epsilon_{0}>0$ small enough. We choose a sequence of positive numbers $\{\epsilon_{j}\}_{j\in\mathbb{N}}$ such that $\epsilon_{j}=2^{-(j+1)}\epsilon_{0}$. Then we define, for $j\in\mathbb{N}$ large enough, $D_{\epsilon_{j}}\coloneqq\{\psi<-\epsilon_{j}\}$, and
$h^{j}_{w}(z)\coloneqq\begin{cases}\breve{f}_{w}(z-\frac{\epsilon_{0}}{2^{j+3}}\hat{N}),\ z\in\overline{D_{\epsilon_{j}}\cap M},\\ \breve{f}_{w}(z),\ z\in M\setminus\bar{D},\\ 0\ \text{otherwise.}\end{cases}$
Let $\{\chi_{j}\}$ be a sequence of smooth cutoff functions where for each $j$, $1-\chi_{j}$ is supported in $U$. Moreover we set $\chi_{j}$ to be identically 1 in $\{\psi\geq\epsilon_{j}\}\cup\{\psi\leq-2\epsilon_{j}\}$ and 00 in $\{-\epsilon_{j}\leq\psi\leq\epsilon_{j}/2\}$. From here we define a sequence of $(n,0)$ forms as $\tilde{f}_{w,j}(z)\coloneqq h^{j}_{w}(z)\chi_{j}(z)dz$ that are smooth up to $\bar{M}$. It is easy to see that $\tilde{f}_{w,j}\to\tilde{f}_{w}$ in $L^{2}_{n,0}(M)$. Set $\tilde{g}_{w,j}(z)\coloneqq\bar{\partial}\tilde{f}_{w,j}.$ We note that $supp(\tilde{g}_{w}(z))\subset\partial D\setminus\partial M$, and by our construction, $supp(\tilde{g}_{w,j}(z))$ is contained in $\{-2\epsilon_{j}<\psi(z)<\epsilon_{j}\}\cap M$ for $j$ large enough. This implies that the Euclidean distance between boundary of $supp(\tilde{g}_{w,j}(z))$ from $supp(\tilde{g}_{w}(z))$ is at most $2^{-j}\epsilon_{0}$, for $j$ large enough.
Moreover, we have
@@
Let $N:L^{2}_{n,1}(M)\to L^{2}_{n,1}(M)$ be the $\bar{\partial}$-Neumann operator. Let the point $p\in\partial M$ have finite type $m$ and $g=\bar{\partial}f$, for a smooth $(n,0)$ form $f$, be compactly supported in $U\cap\overline{M}$. Then the following pseudolocal estimate holds:
@@
for any $\xi,\xi_{1}\in C^{\infty}_{c}(B(p,\frac{3}{2}\delta))$ that satisfy $\xi_{1}|_{W}\equiv 1$ and $\xi|_{B(p,\delta)}\equiv 1$, where $W\subset B(p,\frac{3}{2}\delta)$ is a small neighborhood of $supp(\xi)$. The constants $\{C_{s}\}$ depend on $s$ and $\xi$, but not on $W$ and $\xi_{1}$. In the two dimensional case, $\lambda$ can be taken as $2^{-m}$ (by [Koh72]) and $\leq 1/m$ in higher dimensions (by [Cat87, Theorem 3.3]).
@@
These pseudolocal estimates essentially follow from the subelliptic estimates of the $\bar{\partial}$-Neumann operator, as mentioned in [FK72]. In two dimensions, subelliptic estimates are proven in [Koh72]. In higher dimensions, subelliptic estimates for D’Angelo finite type domains in $\mathbb{C}^{n}$ were proved by Catlin (see [CD10], [Cat87]); these are also valid in our manifold setting because of their local nature. In the following we give a proof of how these pseudolocal estimates follow from the subelliptic ones for manifolds. We also note that the proof of existence of the operator $N$ for $(n,1)$ forms is similar to that for $(0,1)$ forms.
@@
To apply Lemma B.1, we let $\xi,\xi_{1}$ be as in Claim 2.1. Let $d:=dist(supp(\xi),supp(1-\xi_{1}))$ (where the distance is measured with respect to the chart $U$). We note, because of Lemma 2.1, $d<\delta<1$. Furthermore, we take $\epsilon=\lambda$, $r=s+\lambda$, $m=1$, $m^{\prime}=1+2\lambda$, $L=N$, $f=\tilde{g}_{w,j}$, $R(\cdot)=|\!|{}\cdot|\!|{}_{s}$, $T=id$ and $G(\cdot)=|\!|{}\cdot|\!|{}^{\dagger}_{-1}$ in Lemma B.1. By (2.7), we see that the inequality (B.2) holds. Note that to apply Lemma B.1, we also need $N$ to be a linear operator from $L^{2}(M)$ to itself, and that it satisfies (B.1) with $\Omega$ replaced by $M$. This is verified in Proposition C.1 in Appendix C. Applying Lemma B.1 then gives us (2.6). ∎
@@
Let $\sigma>0$ be small and $\rho\in C^{\infty}(M^{\prime})$ be a real-valued function so that $\rho$ is compactly supported in a $2\sigma$ neighborhood of $\overline{\partial D\setminus\partial M}$ in $M^{\prime}$, with respect to the Euclidean metric on $U$, which we denote by $V_{2\sigma}$ and is identically one in $V_{\sigma}$. Write $K_{D}(z,w)=\tilde{K}_{D}(z,\bar{w})\wedge d\bar{w}|_{w}$ for all $w\in B(p,\delta)\cap M$. Fix any $\phi=\sum_{j=1}^{n}\phi_{j}dz_{1}\wedge\cdots\wedge dz_{n}\wedge d\bar{z}_{j}\in\Omega^{n,1}(\overline{M})\cap Dom(\bar{\partial}^{*})$. Since $supp(\tilde{g}_{w})\subset\partial D\setminus\partial M$, by our definition of $\tilde{g}_{w}$ acting on $\Omega^{n,1}(\overline{M})\cap Dom(\bar{\partial}^{*})$, we have $(\tilde{g}_{w},\phi)=(\tilde{g}_{w},\rho\phi)$ and
@@
where $\tilde{K}_{D}(z,\bar{w})=k_{D}(z,\bar{w})dz_{1}\wedge\cdots\wedge dz_{n}$ in local coordinates. Assuming $\sigma$ is sufficiently small, we observe that $dist(V_{2\sigma},B(p,\delta))>0$. By Theorem 1 in [Bel86], it follows that